### v2

@@ 4.1.2 Segmentation
DiS (Fei et al, 2024) Vim Natural Images Image Generation BD (H) Raster
ZigMa (Hu et al, 2024) Mamba* Natural Images Image Generation BD (H/V) Zigzag
DiM (Teng et al, 2024) Mamba Natural Images Image Generation BD (H/V) Raster
DiM (Mo and Tian, 2024) Mamba Natural Images Image Generation BD (H) Raster
Gamba (Shen et al, 2024) Mamba* Natural Images 3D Reconstruction SD (H) Raster
MD-Dose (Fu et al, 2024b) Mamba 2D Medical Images (CT) Radiation Dose Prediction SD (H) Raster
VM-DDPM (Ju and Zhou, 2024) VSS* 2D Medical Images (CXR/MRI) 2D Medical Image Synthesis BD (H/V) Raster
SMCD (Qian et al, 2024) Mamba* Motion Sequence (Style/Content) Motion Style Transfer SD (H) Raster
@@
UVMNet (Zheng and Wu, 2024) Mamba* Natural Images Dehazing/Low Light Enhancement/Deraining SD (H) + Channel Raster
CU-Mamba (Deng and Gu, 2024) Mamba* Natural Images Denoising/Deblurring SD (H) + Channel Raster
ALGNet (Gao and Dang, 2024) Mamba Natural Images Deblurring SD (H) Raster
EVSSM (Kong et al, 2024) VSS* Natural Images Deblurring BD (H/V) Raster
FreqMamba (Zhen et al, 2024) VSS* Natural Images Deraining BD (H/V) + FT + Local Sampling Raster
DFSSM (Yamashita and Ikehara, 2024) VSS* Natural Images Deraining BD (H/V) Raster
FourierMamba (Li et al, 2024a) VSS* Natural Images Deraining BD (H/V) + FT Raster
MambaIR (Guo et al, 2024a) VSS* Natural Images Super-resolution/Denoising BD (H/V) Raster
VmambaIR (Shi et al, 2024b) Mamba* Natural Images Super-resolution/Deraining BD (H/V/D) Raster
MMA (Cheng et al, 2024) Vim* Natural Images Super-resolution BD (H) Raster
DVMSR (Lei et al, 2024) Mamba Natural Images Super-resolution SD (H) Raster
IRSRMamba (Huang et al, 2024g) VSS* Natural Images Super-resolution BD (H/V) Raster
MambaUIE (Chen and Ge, 2024) VSS* Natural Images Underwater Image Enhancement BD (H/V) Raster
WaterMamba (Guan et al, 2024) VSS* Natural Images Underwater Image Enhancement BD (H/V) + Channel Raster
Retinexmamba (Bai et al, 2024) VSS* Natural Images Low Light Enhancement BD (H/V) Raster
FDVMNet (Zheng and Zhang, 2024) Mamba* 2D Medical Images (Endo) Endoscopic Exposure Correction SD (H) Raster
MambaMIR (Huang et al, 2024b, c) VSS* 3D Medical Images (FMRI/SVCT) Medical Image Reconstruction BD (H/V) Raster
MSDR (Li et al, 2024i) VSS* 2D Medical Images (ODT) Sparse Reconstruction BD (H/V) Raster
HSIDMamba (Liu et al, 2024h) VSS*/Vim* Remote Sensing Images Hyperspectral Denoising BD (H/V/H*/V*) Zigzag
SSUMamba (Fu et al, 2024a) Vim* Remote Sensing Images Hyperspectral Denoising BD (H/V/D) Raster
RSDehamba (Zhou et al, 2024a) VSS* Remote Sensing Images Dehazing BD (H/V) Raster
FMSR (Xiao et al, 2024) VSS* Remote Sensing Images Super-resolution BD (H/V) Raster
GMSR (Wang et al, 2024e) VSS* Remote Sensing Images Spectral Reconstruction BD (H/V) Raster
@@
MambaAD (He et al, 2024a) VSS* Natural Images Multi-class Anomaly DET BD (H/V/H*/V*) Hilbert
MambaVC (Qin et al, 2024) VSS Natural Images Visual Compression BD (H/V) Raster
4.1.3 Detection
@@
FER-YOLO-Mamba (Ma et al, 2024a) is the first Mamba-based model for facial expression detection. It introduced a dual-branch structure that incorporated omnidirectional Mamba and attention mechanisms for the Feature Pyramid Network (FPN). VMambaCC (Ma et al, 2024c) is a pioneering work that applied the Mamba architecture to crowd counting tasks. It used VSS blocks for feature extraction and combined VSS blocks with attention mechanisms for feature fusion. SOAR (Verma et al, 2024) integrated vision Mamba block, which utilizes position embedding and bidirectional Mamba for effective visual modeling, into the YOLO v9 architecture. This integration enhances the real-time detection performance of small aerial objects.
4.1.4 Generation
@@
Intuitively, applying the Mamba architecture to a series of generation tasks to achieve sufficient long sequence interactions has the potential to achieve impressive performance.
DiS (Fei et al, 2024) leverages a state space-based architectural backbone and treats all inputs, including time, condition and noisy image patches, as tokens, which enables the model to effectively capture long-range dependencies within the data. It achieves compared performance in both unconditional and class-conditional image generation scenarios. ZigMa (Hu et al, 2024) and DiM (Mo and Tian, 2024) are DiT-style (Peebles and Xie, 2023) Mamba diffusion models. Additionally, ZigMa (Hu et al, 2024) leveraged spatial continuity to maximally incorporate the inductive bias from visual data. DiM (Teng et al, 2024) was another diffusion Mamba model specifically designed for efficient high-resolution image synthesis. The authors implemented multi-directional scans to adapt Mamba for 2D signal processing. Furthermore, they introduced learnable padding tokens at the end of each row and column to enhance the model’s sensitivity to discontinuities between rows or columns, along with lightweight local feature enhancement. Gamba (Shen et al, 2024) introduced GambaFormer to process 3D Gaussian Splatting by utilizing the newly proposed Gaussian Mamba.
MD-Dose (Fu et al, 2024b) introduced a novel diffusion model based on the Mamba architecture for predicting radiation therapy dose distribution in thoracic cancer patients. VM-DDPM (Ju and Zhou, 2024) is the first to leverage an SSM-CNN hybrid architecture in diffusion models for medical image synthesis. SMCD (Qian et al, 2024) is the first style motion conditional diffusion framework that treats style motion as a condition. It is also the first to apply Mamba to predict stylized motion in each sampling step.
4.1.5 Restoration
@@
Recently, the Mamba architecture has also been extensively applied to various image restoration tasks. These low-level tasks include reconstructing degraded images to their original state, such as image dehazing (Zheng and Wu, 2024), deraining (Zheng and Wu, 2024; Zhen et al, 2024), and super-resolution (Guo et al, 2024a; Cheng et al, 2024; Lei et al, 2024), and enhancing the image quality by adjusting luminance, contrast, and other visual attributes, such as low light enhancement (Zheng and Wu, 2024; Bai et al, 2024), and underwater image enhancement (Chen and Ge, 2024).
The extraction of local fine-grained features plays a critical role in tackling low-level image tasks such as dehazing, which the vanilla Mamba architecture lacks. Incorporating convolutions can enhance the visual representation of Mamba-based methods (Zheng and Wu, 2024). Additionally, the correlation between channels is vital. UVMNet (Zheng and Wu, 2024) and CU-Mamba (Deng and Gu, 2024) processed the feature maps in both spatial and channel domains, in parallel and serially respectively. In image deblurring tasks, ALGNet (Gao and Dang, 2024) utilized simplified channel attention to model local connectivity. EVSSM (Kong et al, 2024) applied various geometric transformations before each SSM-based module instead of the common practice of scanning in four directions for efficient computation. It also introduced a frequency-domain-based Feedforward network. Image restoration approaches often utilize techniques from the frequency domain, leveraging the ability to isolate and manipulate specific image frequency components for improving detail recovery. For instance, high-frequency rain streaks in draining tasks are removed using frequency techniques (Zhen et al, 2024). Additionally, DFSSM (Yamashita and Ikehara, 2024) incorporated a mixed-scale gated convolutional block, and FourierMamba (Li et al, 2024a) employed a zigzag scanning strategy specifically designed for the frequency domain. Similarly, in image super-resolution tasks, techniques such as convolutions (Guo et al, 2024a; Lei et al, 2024), channel attention (Guo et al, 2024a; Cheng et al, 2024), channel scanning (Shi et al, 2024b), and frequency analysis (Huang et al, 2024g) are applied. Additionally, DVMSR (Lei et al, 2024) employed a distillation strategy to enhance both efficiency and performance. In image enhancement tasks, MambaUIE (Chen and Ge, 2024) and WaterMamba (Guan et al, 2024) were specifically developed to improve the visual quality of underwater images degraded due to light absorption and scattering, integrating convolutions and spatial-channel scanning associate with multi-scale Feedforward network respectively. Retinexmamba (Bai et al, 2024) incorporated illumination-fused attention and the Mamba architecture for low light enhancement.
In the medical domain, FDVMNet (Zheng and Zhang, 2024) introduced a C-SSM block by combining convolution, SSM, cross-attention, and shortcut as a basic cell, and utilized two paths to process the phase and amplitude of the frequency domain information respectively. MambaMIR (Huang et al, 2024b, c) introduced an Arbitrary-Masked State Space (AMSS) block based on VSS by adopting an arbitrary-mask mechanism to leverage the scanning redundancy. MSDR (Li et al, 2024i) combined the VSS block and an Inception-based FeedForward Network module.
In the remote sensing domain, denoising and dehazing methods include zigzag scanning in eight spatial directions (Liu et al, 2024h), Spatial-Spectral Alternating Zigzag Scan strategy (Fu et al, 2024a), and Direction-aware Scan Module (Zhou et al, 2024a). Additionally, spectral attention (Liu et al, 2024h) and channel attention (Zhou et al, 2024a) were employed. FMSR (Xiao et al, 2024) was developed to improve spatial-frequency fusion in remote sensing image super-resolution tasks. GMSR (Wang et al, 2024e) introduced a gradient-guided Mamba for spectral reconstruction from RGB images.
4.1.6 Others
Some works employ Mamba for other vision tasks. MambaAD (He et al, 2024a) applied Mamba to multi-class unsupervised anomaly detection. MambaVC (Qin et al, 2024) explored the use of Mamba architecture for visual compression.
Table 6: The Mamba-based methods in video-level vision tasks. The abbreviations of Block here are VSS: Vision State Space Module, Vim: Vision Mamba, *: modification. The abbreviations of Task here are CLS: Classification, SEG: Segmentation
Methods Block Data Task Code
@@ 4.1.6 Others
Vivim (Yang et al, 2024e) Vim* Ultrasound Videos, Endoscopic Videos Video Object SEG
VideoMamba (Li et al, 2024c) Vim* Natural Videos Action CLS, Long-form Video Understanding, Text-to-Video Retrieval
VideoMamba Suite (Chen et al, 2024a) Vim* Natural Videos Action Localization, Action anticipation, Action Recognition, Video Captioning
Video Grounding, Video Question Answering, Multi-instance Retrieval
RhythmMamba (Zou et al, 2024a) Mamba* Facial Videos rPPG Prediction
Simba (Chaudhuri and Bhattacharya, 2024) Mamba Skeletal Action Videos Skeletal Action Recognition
DeMamba (Chen et al, 2024b) Mamba Natural Videos AI-Generated Video Detection
MAMBA4D (Liu et al, 2024b) Mamba Point Cloud Videos 3D Action Recognition, 4D Semantic SEG
@@
SSMDiff (Oshima et al, 2024) Vim Natural Videos Video Generation
@@
ZigMa (Hu et al, 2024) Mamba* Natural Videos Video Generation
Matten (Gao et al, 2024) Vim* Natural Videos Video Generation
DiM (Mo and Tian, 2024) Mamba Natural Videos Video Generation
4.2 Video

### v3

@@ 4.1.1. Classification
ZigMa (Hu et al., 2024) Natural Images Image Generation 1D BD H/V Zigzag Global ECCV24
VSSM (Lambrechts et al., 2024) Natural Images Image Generation 1D SD H Raster Global ICML24 Workshop
@@
FreqMamba (Zhen et al., 2024) Natural Images Deraining 2D BD H/V + F Raster Global/Local MM24
MambaIR (Guo et al., 2024a) Natural Images Super-resolution/Denoising 2D BD H/V Raster Global ECCV24
DVMSR (Lei et al., 2024) Natural Images Super-resolution 1D SD H Raster Global CVPR24 Workshop
MambaLLIE (Weng et al., 2024) Natural Images Enhancement 2D BD H/V Raster Global NeurIPS24
ECMamba (Dong et al., 2024c) Natural Images Enhancement 2D BD H/V Reordering Global NeurIPS24
Deform-Mamba (Ji et al., 2024) 2D Medical Images (MRI) Super-resolution 2D BD H/V Raster Global MICCAI24
SSUMamba (Fu et al., 2024c) Remote Sensing Images Hyperspectral Denoising 1D BD H/V/D Raster Global TGRS24
FMSR (Xiao et al., 2024c) Remote Sensing Images Super-resolution 2D BD H/V Raster Global TMM24
MambaFormerSR (Zhi et al., 2024) Remote Sensing Images Super-resolution 2D BD H/V Raster Global GRSL24
VMambaSCI (Zhang et al., 2024c) Remote Sensing Images Spectral Reconstruction 2D BD H/V + C Raster Global MM24
@@
MambaAD (He et al., 2024a) Natural Images Multi-class Anomaly DET 2D BD H/V/H*/V* Hilbert Global NeurIPS24
MTMamba (Lin et al., 2024) Natural Images Multi-task Dense SU 2D BD H/V Raster Global ECCV24
Hamba (Dong et al., 2024a) Natural Images 3D Hand Reconstruction 2D BD Graph-guided Raster Global NeurIPS24
SUM (Hosseini et al., 2025) Natural Images Visual Attention Modeling 2D BD H/V Raster Global WACV25
4.1.2. Segmentation
Segmentation remains a vital and prominent area in computer vision, holding immense value for diverse real-world applications (Chen et al., 2024f; Du et al., 2024b; Ma et al., 2024b; Xing et al., 2024; Ruan and Xiang, 2024; Liu et al., 2024g; Gong et al., 2024; Wang et al., 2024o; Ma and Wang, 2024; Ye and Chen, 2024; Wang and Ma, 2024; Liao et al., 2024; Wang et al., 2024a; Zhang et al., 2024h; Wu et al., 2024a; Wang et al., 2024b; Xie et al., 2024b; Tang et al., 2024a; Sanjid et al., 2024a; Wu et al., 2024b; Hao et al., 2024; Archit and Pape, 2024; Isensee et al., 2024; Sanjid et al., 2024b; Nguyen et al., 2024; Xu, 2024; Yuan et al., 2024; Tsai et al., 2024; Chen et al., 2024d; Shi et al., 2024c; Wang et al., 2024g; Chang et al., 2024; Fan et al., 2024; Zhou et al., 2024d; Zhu et al., 2024a; Ma et al., 2024d; Yao et al., 2024; Wang et al., 2024l; Huang et al., 2024a; Zhu et al., 2024b; Liu et al., 2024b; Zhao et al., 2024a; Chen et al., 2024e; Zhou et al., 2024a; He et al., 2024c; Li et al., 2024e; Malik et al., 2024; Wei et al., 2024; Khan et al., 2025). Mamba is well-positioned to enhance segmentation tasks for its strong long-range modeling capability.
In the context of few-shot segmentation, HMNet (Xu et al., 2024b) identified two issues when concatenating support and query features and applying the original Mamba: support forgetting and the intra-class gap between support and query. To mitigate these issues, HMNet introduced two mechanisms: periodically re-scanning the support features during the query feature scanning process, and scanning the query features in parallel to avoid undesired pixel-level interactions within the query features.
Recent advances based on Mamba have appeared in the medical domain, encompassing diverse 2D and 3D medical image datasets (Tsai et al., 2024; Chen et al., 2024d). Several methods (Archit and Pape, 2024; Liu et al., 2024g; Zhang et al., 2024h) directly replace CNN-blocks in the U-Net architecture with Mamba-based blocks (Gu and Dao, 2024; Liu et al., 2024e; Zhu et al., 2024c). U-Mamba (Ma et al., 2024b) is the first effective attempt to apply Mamba in medical image segmentation. It introduced a hybrid CNN-SSM block. ViM-UNet (Archit and Pape, 2024) utilized Vim (Zhu et al., 2024c) block as the basic blocks. Swin-UMamba (Liu et al., 2024g) and VM-UNet-V2 (Zhang et al., 2024h) employed the VMamba (Liu et al., 2024e) block as a basic block to capture extensive contextual information. MM-UNet (Du et al., 2024b), Polyp-Mamba (Xu et al., 2024c), GGVMamba (Zhou et al., 2024d), and LKM-UNet (Wang et al., 2024b) employed Mamba-based architecture as a basic block with additional designs for enhancement. SegMamba (Xing et al., 2024) introduced gated spatial convolution and tri-orientated Spatial Mamba, which utilized an inter-slice path for 3D data. EM-Net (Chang et al., 2024) employed a channel scanning strategy and integrated frequency analysis. CAMS-Net (Khan et al., 2025) introduced Mamba-based channel and spatial aggregators. TP-Mamba (Wang et al., 2024g) and ShapeMamba-EM (Shi et al., 2024c) each introduced a 3D adapter based on Mamba to fine-tune the foundation models SAM (Kirillov et al., 2023) and SAM-Med3D (Wang et al., 2024d), respectively, for medical image data. PathMamba (Fan et al., 2024) incorporated Mamba into weakly supervised multi-label segmentation of pathology images and introduced a dual-granularity comparative Mamba block, designed to facilitate pixel- and patch-level contrastive learning. In (Ma and Wang, 2024), a purely Mamba-based UNet and a CNN-based UNet were integrated into a semi-supervised learning framework, where they perform cross-supervision and contrastive learning.
In the field of remote sensing images, Mamba-based approaches typically leverage its efficiency for processing high-resolution images (Ma et al., 2024d), and exploit the spatial-temporal and spatial-spectral characteristics of images to perform scanning (Chen et al., 2024e; Li et al., 2024e).
4.1.3. Detection
@@
FER-YOLO-Mamba (Ma et al., 2024a) is the first Mamba-based model for facial expression detection. It introduced a dual-branch structure that incorporated omnidirectional Mamba and attention mechanisms for the feature pyramid network. VMambaCC (Ma et al., 2024e) is a pioneering work that applied the VMamba (Liu et al., 2024e) architecture to crowd counting tasks. SOAR (Verma et al., 2024) integrated vim (Zhu et al., 2024c) block into the YOLO v9 architecture.
4.1.4. Generation
@@
Intuitively, applying the Mamba architecture to a series of generation tasks to achieve sufficient long sequence interactions has the potential to achieve impressive performance (Fei et al., 2024; Hu et al., 2024; Teng et al., 2024; Mo and Tian, 2024; Lambrechts et al., 2024; Shen et al., 2024; Fu et al., 2024a; Ju and Zhou, 2024; Qian et al., 2024).
ZigMa (Hu et al., 2024) is a DiT-style (Peebles and Xie, 2023) Mamba diffusion model. It leveraged spatial continuity to maximally incorporate the inductive bias from visual data. VSSM (Lambrechts et al., 2024) is a variational autoencoder in which both the encoder and decoder are SSMs, enabling efficient parallel computation.
4.1.5. Restoration
@@
Recently, the Mamba architecture has also been extensively applied to various image restoration tasks. These low-level tasks include reconstructing degraded images to their original state, such as image dehazing (Zheng and Wu, 2024; Zhou et al., 2024b), deraining (Zheng and Wu, 2024; Zhen et al., 2024; Yamashita and Ikehara, 2024; Li et al., 2024d; Shi et al., 2024d), denoising (Deng and Gu, 2024; Guo et al., 2024a; Liu et al., 2024f; Fu et al., 2024c), deblurring (Deng and Gu, 2024; Gao and Dang, 2024; Kong et al., 2024), super-resolution (Guo et al., 2024a; Shi et al., 2024d; Cheng et al., 2024; Lei et al., 2024; Huang et al., 2024c; Ji et al., 2024; Xiao et al., 2024c; Zhi et al., 2024), and other reconstruction (Zheng and Zhang, 2024; Huang et al., 2024g, h; Li et al., 2024g; Wang et al., 2024e; Zhang et al., 2024c), and enhancing the image quality by adjusting luminance, contrast, and other visual attributes, such as multiple exposure correction (Tan et al., 2024), low light enhancement (Zheng and Wu, 2024; Weng et al., 2024; Bai et al., 2024), and underwater image enhancement (Chen and Ge, 2024; Guan et al., 2024).
The extraction of local fine-grained features plays a critical role in tackling low-level image tasks, which the vanilla Mamba architecture lacks. Besides, image restoration approaches often utilize techniques from the frequency domain, leveraging the ability to isolate and manipulate specific image frequency components for improving detail recovery. For instance, FreqMamba (Zhen et al., 2024) utilized both Global and Local scanning, effectively removing high-frequency rain streaks in deraining tasks by leveraging frequency-based techniques. Similarly, in image super-resolution tasks, MambaIR (Guo et al., 2024a) incorporated convolutions and channel attentions to enhance the Mamba architecture’s capacity for preserving local information while effectively reducing channel redundancy. Differently, DVMSR (Lei et al., 2024) applied a distillation strategy to the Mamba-based network to enhance its efficiency. In image enhancement, MambaLLIE (Weng et al., 2024) enhanced local dependencies by incorporating an augmented local bias within the 2D selective scan mechanism and introducing an implicit Retinex-aware selective kernel module. ECMamba (Dong et al., 2024c) integrated Retinex theory for exposure correction and presented a feature-aware scanning strategy that leverages an activation response map from deformable convolution to generate sequences ordered by feature importance.
In the medical domain, Deform-Mamba (Ji et al., 2024) combined the modulated deformable block with VMamba (Liu et al., 2024e) block to effectively capture both local and global information. Additionally, it introduced a multi-view context module and a contrastive edge loss.
In the remote sensing domain, to improve computational efficiency in hyperspectral denoising, SSUMamba (Fu et al., 2024c) introduced a spatial-spectral continuous scan Mamba, alternating rows, columns, and bands in six different orders to capture long-range spatial-spectral dependencies. Furthermore, 3D convolutions are incorporated to enhance local spatial-spectral modeling. In remote sensing image super-resolution tasks, FMSR (Xiao et al., 2024c) enhanced spatial-frequency fusion by integrating frequency selection modules with the global VMamba (Liu et al., 2024e) block and a local hybrid gating module. MambaFormerSR (Zhi et al., 2024) combined Mamba with Transformer architectures, while also introducing a convolutional Fourier transform feed-forward network. VMambaSCI (Zhang et al., 2024c) introduced a novel spatial-channel scanning technique for spectral reconstruction.
4.1.6. Others
Mamba’s ability to model long-range dependencies extends its advantages to various other image modality tasks (He et al., 2024a; Qin et al., 2024; Lin et al., 2024; Dong et al., 2024a; Hosseini et al., 2025), improving feature representation and contextual understanding across diverse applications. MambaAD (He et al., 2024a) applied Mamba to multi-class unsupervised anomaly detection by combining a pre-trained encoder with a Mamba-based decoder. The decoder integrated parallel cascaded hybrid state space blocks, utilizing Hilbert scanning in eight directions, and multi-kernel convolution operations at multiple scales to effectively capture both global and local information. MTMamba (Lin et al., 2024) introduced self-task and cross-task Mamba blocks for multi-task dense prediction. In particular, the cross-task Mamba block was designed to enhance interactions between tasks. Hamba (Dong et al., 2024a) introduced a graph-guided bidirectional scanning technique for 3D hand reconstruction from a single RGB image, efficiently utilizing a few effective tokens to learn the joint relations and spatial sequences. SUM (Hosseini et al., 2025) leveraged Mamba for efficient long-range dependency modeling and introduced a novel conditional visual state space block, enabling dynamic adaptation to different image types, in visual attention modeling tasks.
Table 7. The Mamba-based methods in video modality vision tasks. The abbreviations of Task here are CLS: Classification, VU: Video Understanding, T2VR: Text-to-Video Retrieval, MOT: Multi-object Tracking, STF: Spatiotemporal Forecasting, VFI: Video Frame Interpolation.
Methods Data Task Scan Strategy Venue Code
@@ 4.1.6. Others
VideoMamba (Li et al., 2024c) Natural Videos Action CLS, VU, T2VR BD Spatial-first ECCV24
VideoMamba (Park et al., 2024a) Natural Videos VU BD Spatial-first ECCV24
MambaTrack (Xiao et al., 2024a) Natural Videos MOT BD Spatial-first MM24 Oral
DLT (Jin et al., 2024) Natural Videos MOT BD Spatial-first MM24
VMRNN (Tang et al., 2024b) Natural Videos STF Spatial CVPR24 Workshop
@@
RainMamba (Wu et al., 2024c) Natural Videos Video Deraining BD Spatial-first (Zigzag) + 3D Hilbert MM24 Oral
@@
SSMDiff (Oshima et al., 2024) Natural Videos Video Generation Spatial + Temporal ICLR24 Workshop
@@
ZigMa (Hu et al., 2024) Natural Videos Video Generation Spatial + Temporal ECCV24
VFIMamba (Zhang et al., 2024e) Natural Videos VFI Temporal-first NeurIPS24
MMD (Dong et al., 2024b) Motion Sequence Motion Generation Spatial + Temporal MM24
Figure 7. Video scanning techniques.
[…]
4.2. Video